# Positive and nodal limiting profiles for a semilinear elliptic equation with a shrinking region of attraction

### v1

@@ 1 Introduction
where $\varepsilon>0$, $Q_{\varepsilon}(x):=Q(\frac{x}{\varepsilon})$, i.e., $Q_{\varepsilon}(x)=1$ if $x\in\varepsilon\Omega$ and $Q_{\varepsilon}(x)=-1$ if $x\in\mathbb{R}^{N}\smallsetminus\varepsilon\Omega$, $p\in(2,2^{*})$ and $2^{*}$ is the critical Sobolev exponent, namely, $2^{*}:=\frac{2N}{N-2}$ if $N\geq 3$ and $2^{*}=\infty$ if $N=1,2$.
@@
where $E:=D^{1,2}(\mathbb{R}^{N})\cap L^{p}(\mathbb{R}^{N})$. Thus, it is of interest to obtain information on the qualitative properties of the solutions to (1.2). This is one of our goals.
@@
The following result is an easy extension of [7, Theorem 1.1] to an arbitrary domain $\Omega$.
@@
has only the trivial solution for any $\kappa\in\mathbb{R}$. This is obvious if $\kappa\leq 0$ and it is a consequence of the Pohozhaev identity [18, Theorem B.3] if $\kappa>0$.
@@
    Let $N\geq 2$ and let $\Omega$ be a radially symmetric open bounded subset of $\mathbb{R}^{N}$ with smooth boundary. Then, any least energy solution and any least energy nodal solution of (1.2) is foliated Schwarz symmetric in $\mathbb{R}^{N}$.
@@
Finally, we study the decay at infinity of solutions to (1.2) in dimensions $N\geq 3$. We write $B_{\varrho}$ for the ball of radius $\varrho$ in $\mathbb{R}^{N}$ centered at the origin.
@@
    $w(x)\geq C_{\delta}|x|^{2-N+\delta}\qquad\text{for every \ }x\in\mathbb{R}^{N}\smallsetminus B_{\varrho}.$
@@
Establish the precise decay of the solutions to problem (1.2) for every $p\in(2,2^{*})$ and $N\geq 1$.
@@ 2 Existence of positive least energy solutions
Let $u\in H^{1}(\mathbb{R}^{N})$ be a solution of (2.1). Setting $a(x)=\varepsilon^{2}+|u(x)|^{p-2}$ we see that $|-\varepsilon^{2}u+Q(x)|u|^{p-2}u|\leq a(x)|u|$ and $a\in L^{\frac{N}{2}}_{loc}(\mathbb{R}^{N})$. Then, a Brezis-Kato-Moser argument shows that $u\in L^{s}_{loc}(\mathbb{R}^{N})$ for every $s\in[1,\infty)$ (see [13, Lemma B.3]), and standard elliptic regularity arguments yield that $u\in\mathcal{C}^{1,\alpha}_{loc}(\mathbb{R}^{N})$ for every $\alpha\in(0,1)$ and $u\in\mathcal{C}_{loc}^{\infty}(\Omega)\cap\mathcal{C}_{loc}^{\infty}(\mathbb{R}^{N}\smallsetminus\overline{\Omega})$ (see [13, Appendix B]). ∎
@@
The existence of a least energy solution follows from Ekeland’s variational principle and Lemma 2.3. Replacing it by its absolute value, we obtain a least energy solution such that $u\geq 0$ in $\mathbb{R}^{N}$. As shown in Lemma 2.4, $u\in W^{2,N}_{loc}(\mathbb{R}^{N})$. Thus, the strong maximum principle for strong solutions [8, Theorem 9.6] applied to the equation
@@
is continuous and odd. Therefore, $\infty=\mathrm{genus}(\Sigma_{\Omega})\leq\mathrm{genus}(\mathcal{N}_{\varepsilon})$, where “$\mathrm{genus}$” stands for the Kransnoselskii genus. It follows from [13, Theorem II.5.7] that $J_{\varepsilon}$ has infinitely many pairs of critical points on $\mathcal{N}_{\varepsilon}$ and following the argument in [11, Proposition 9.33] one shows that $J_{\varepsilon}$ has an unbounded sequence of critical values on $\mathcal{N}_{\varepsilon}$, as claimed. ∎
@@ 3 The limit profile of positive least energy solutions
When $\Omega$ is the unit ball, the behavior of positive least energy solutions to (2.1) as $\varepsilon\to 0$ is described in [7, Theorem 1.1]. It is shown that, after passing to a subsequence they converge strongly in $E$ to a positive least energy solution to the problem (1.2). The general case is similar. We give the details.
@@
that $u_{k}\to u$ strongly in $D^{1,2}(\mathbb{R}^{N})$ and that $t=1$. Hence, $\|u\|^{2}=\int_{\mathbb{R}^{N}}Q(x)|u|^{p}=\frac{2p}{p-2}c_{0}$. Combining this identity with (3.4) we see that $u_{k}\to u$ strongly in $L^{p}(\mathbb{R}^{N})$ and, thus, in $E$. As $u\in\mathcal{N}_{0}$ and $J_{0}(u)=c_{0}$, $u$ is a least energy solution to (1.2). Since $u\geq 0$, from Lemma 3.3 and the maximum principle (see, for instance, [8, Theorem 9.6]) we derive that $u>0$ in $\mathbb{R}^{N}$. ∎
@@ 4 Existence of least energy nodal solutions
$\gamma(s)\in\widehat{\mathcal{N}}_{\varepsilon}^{+}\text{ \ if \ }s\in(0,\tfrac{1}{2})\qquad\text{and}\qquad\gamma(s)\in\widehat{\mathcal{N}}_{\varepsilon}^{-}\text{ \ if \ }s\in(\tfrac{1}{2},1).$
@@
Given $w\in\widehat{\mathcal{N}}_{\varepsilon}$ and $s\in[0,1]$, let $t_{w^{+}},t_{w^{-}}\in(0,\infty)$ be such that $t_{w^{+}}w^{+}\in\mathcal{N}_{\varepsilon}$ and $t_{w^{-}}w^{-}\in\mathcal{N}_{\varepsilon}$, namely,
$t_{w^{+}}=\left(\frac{\|w^{+}\|_{\varepsilon}^{2}}{\int_{\mathbb{R}^{N}}Q(x)|w^{+}|^{p}}\right)^{\frac{1}{p-2}}\quad\text{ and }\quad t_{w^{-}}=\left(\frac{\|w^{-}\|_{\varepsilon}^{2}}{\int_{\mathbb{R}^{N}}Q(x)|w^{-}|^{p}}\right)^{\frac{1}{p-2}}.$
Then, let us define
$w_{s}:=(1-s)t_{w^{+}}w^{+}+st_{w^{-}}w^{-}\in\mathcal{U}$
and, by Lemma 2.1, there exists a unique positive number $t_{w_{s}}$ such that $\gamma_{w}(s):=t_{w_{s}}w_{s}\in\mathcal{N}_{\varepsilon}$. As $w\in\mathcal{N}_{\varepsilon}$, the map $\gamma_{w}$ satisfies
$\gamma_{w}(0)=t_{w^{+}}w^{+},\qquad\gamma_{w}(1)=t_{w^{-}}w^{-},\qquad\gamma_{w}\left(\frac{t_{w^{+}}}{t_{w^{+}}+t_{w^{-}}}\right)=t_{w_{s}}\frac{t_{w^{+}}t_{w^{-}}}{t_{w^{+}}+t_{w^{-}}}w=w.$
Therefore, $t_{w_{s}}=\frac{t_{w^{+}}+t_{w^{-}}}{t_{w^{+}}t_{w^{-}}}$. Note that
$w\in\widehat{\mathcal{N}}_{\varepsilon}^{+}\quad\text{ \ if and only if \ }\quad t_{w^{+}}<1,\ t_{w^{-}}>1\text{ \ and \ }0<\frac{t_{w^{+}}}{t_{w^{+}}+t_{w^{-}}}<\frac{1}{2}.$ (4.1)
Now, let $u\in\mathcal{E}_{\varepsilon}$. Then $t_{u^{+}}=1=t_{u^{-}}$, $\gamma_{u}(0)=u^{+}$, $\gamma_{u}(1)=u^{-}$ and $\gamma_{u}(\tfrac{1}{2})=u$. Fix $s\in(0,\frac{1}{2})$ and set $w:=\gamma_{u}(s)$. Then $w^{+}=2(1-s)u^{+}$, $w^{-}=2su^{-}$ and $w\in\widehat{\mathcal{N}}_{\varepsilon}$. From (4.1) we get that $w\in\widehat{\mathcal{N}}_{\varepsilon}^{+}$. Note that $J_{\varepsilon}(t_{w^{+}}w^{-})=J_{\varepsilon}(\frac{s}{1-s}u^{-})>0$. Then, using Lemma 2.1,
@@
$<J_{\varepsilon}(w)=J_{\varepsilon}(2(1-s)u^{+})+J_{\varepsilon}(2su^{-})<J_{\varepsilon}(u^{+})+J_{\varepsilon}(u^{-})=J_{\varepsilon}(u).$
The argument for $s\in(\frac{1}{2},1)$ is similar. ∎
@@
Arguing by contradiction, assume that $J_{\varepsilon}^{\prime}(u)\neq 0$. Fix $\delta>0$ such that $\overline{B}_{\delta}(u):=\{v\in\mathcal{N}_{\varepsilon}:\|v-u\|_{\varepsilon}\leq\delta\}\subset\widehat{\mathcal{N}}_{\varepsilon}$ and $\|\nabla_{\mathcal{N}_{\varepsilon}}J_{\varepsilon}(v)\|\geq\alpha>0$ for every $v\in\overline{B}_{\delta}(u)$, where $\nabla_{\mathcal{N}_{\varepsilon}}J_{\varepsilon}(v)$ is the gradient of $J_{\varepsilon}(v)|_{\mathcal{N}_{\varepsilon}}$ at $v$. Choose $\beta\in(0,\frac{\delta\alpha}{16})$. The deformation lemma [18, Lemma 5.15] yields a map $\eta:[0,1]\times\mathcal{N}_{\varepsilon}\to\mathcal{N}_{\varepsilon}$ such that $\eta(t,u)=u$ if $t=0$ or $u\notin J_{\varepsilon}^{-1}[\widehat{c}_{\varepsilon}-2\beta,\widehat{c}_{\varepsilon}+2\beta]\cap\overline{B}_{\delta}(u)$, $J_{\varepsilon}(\eta(1,u))\leq\widehat{c}_{\varepsilon}-\beta$ and $t\mapsto J_{\varepsilon}(\eta(t,v))$ is nonincreasing for every $v\in\mathcal{N}_{\varepsilon}$. Furthermore, $\eta(t,v)\in\overline{B}_{\delta}(u)$ if $v\in\overline{B}_{\delta}(u)$ for every $t\in[0,1]$. Define $\sigma:[0,1]\to\mathcal{N}_{\varepsilon}$ by $\sigma(s):=\eta(1,\gamma(s))$. Then,
@@ 6.1 Radial symmetry of the positive least energy solution
We recall the notion of Schwarz symmetrization, also called symmetric rearrangement (see [9, Chapter 1] or [10]). Let $A$ be a measurable set of finite volume $|A|$ in $\mathbb{R}^{N}$. Its Schwarz symmetrization $A^{*}$ is the open ball in $\mathbb{R}^{N}$ centered at the origin whose volume is the same as the volume of $A$. Let $u:\mathbb{R}^{N}\to\mathbb{R}$ be a nonnegative measurable function that vanishes at infinity, in the sense that its level sets have finite volume, i.e.,
@@
These and other properties may be found in [10, Section 3.3].
@@
Let $w$ be a positive least energy solution of (1.2) with $\Omega=B_{1}$. By [10, Lemma 7.17] we have that $\|w\|^{2}\geq\|w^{*}\|^{2}$ and, as pointed out above, $|w|_{p}=|w^{*}|_{p}$. We claim that
@@ 6.2 Foliated Schwarz symmetry of the least energy nodal solution
Let $N\geq 2$ and $\mathbb{S}^{N-1}=\left\{x\in\mathbb{R}^{N}:|x|=1\right\}$ be the unit sphere. For each $e\in\mathbb{S}^{N-1}$ we consider the halfspace $H(e):=\{x\in$ $\left.\mathbb{R}^{N}:x\cdot e>0\right\}$, and we denote by $u_{e}$ the composition of the function $u:\mathbb{R}^{N}\rightarrow\mathbb{R}$ with the reflection $\sigma_{e}$ with respect to the hyperplane $\partial H(e)$, that is,
@@
Observe that, for any $u:\mathbb{R}^{N}\to\mathbb{R}$, we have that $(u^{\pm})_{H}=(u_{H})^{\pm}$ and that $(u\chi_{A})_{H}=u_{H}\chi_{A}$ for any radially symmetric set $A$. From [17, Corollary 3.2] we have that
@@
and, from [17, Lemma 3.1], we derive
@@
Let $\Omega$ be a radially symmetric open set in $\mathbb{R}^{N}$ with $N\geq 2.$ There exists $e^{*}\in\mathbb{S}^{N-1}$ such that $u\in\mathcal{C}^{0}(\overline{\Omega})$ is foliated Schwarz symmetric with respect to $e^{*}$ if and only if for every $e\in\mathbb{S}^{N-1}$ either
@@
The proof is given in [3] (see also [12, Proposition 3.2] and [17]). This proof only takes into account the case when $\Omega$ is connected, but the general case easily follows since every component satisfies the same inequalities which completely determine the symmetry axis and the angular monotonicity. ∎
@@
Clearly, $|u-u_{e}|=0$ on $\partial H(e)$. By the maximum principle for strong solutions [8, Theorem 9.6], either $u\equiv u_{e}$ in $H(e)\cap\Omega$ or $\left|u-u_{e}\right|>0$ in $H(e)\cap\Omega$. Since $e$ is arbitrary, Lemma 6.3 states that there exists $e^{*}\in\mathbb{S}^{N-1}$ such that $u$ is foliated Schwarz symmetric in $\Omega$ with respect to $e^{*}$. In particular, it is axially symmetric in $\Omega$ with respect to $\mathbb{R}e^{*}$. Therefore,

### v2

@@ 1 Introduction
where $\varepsilon>0$, $Q_{\varepsilon}(x):=Q(\frac{x}{\varepsilon})$, i.e., $Q_{\varepsilon}(x)=1$ if $x\in\varepsilon\Omega$ and $Q_{\varepsilon}(x)=-1$ if $x\in\mathbb{R}^{N}\smallsetminus\varepsilon\Omega$, $p\in(2,2^{*})$ and $2^{*}$ is the critical Sobolev exponent, namely, $2^{*}:=\frac{2N}{N-2}$.
@@
where $E:=D^{1,2}(\mathbb{R}^{N})\cap L^{p}(\mathbb{R}^{N})$. See also [7] where competitive Schrödinger systems with shrinking regions of attraction are considered, and (1.2) also appears as a limit problem in some cases. Thus, it is of interest to obtain information on the qualitative properties of the solutions to (1.2). This is one of our goals.
@@
The following result is an easy extension of [8, Theorem 1.1] to an arbitrary domain $\Omega$.
@@
has only the trivial solution for any $\kappa\in\mathbb{R}$. This is obvious if $\kappa\leq 0$ and it is a consequence of the Pohozhaev identity [19, Theorem B.3] if $\kappa>0$.
@@
    Let $\Omega$ be a radially symmetric open bounded subset of $\mathbb{R}^{N}$ with smooth boundary. Then, any least energy solution and any least energy nodal solution of (1.2) is foliated Schwarz symmetric in $\mathbb{R}^{N}$.
@@
Finally, we study the decay at infinity of solutions to (1.2). We write $B_{\varrho}$ for the ball of radius $\varrho$ in $\mathbb{R}^{N}$ centered at the origin.
@@
    $w(x)\geq C_{\delta}|x|^{2-N-\delta}\qquad\text{for every \ }x\in\mathbb{R}^{N}\smallsetminus B_{\varrho}.$
@@
Establish the precise decay of the solutions to problem (1.2) for every $p\in(2,2^{*})$.
@@ 2 Existence of positive least energy solutions
Let $u\in H^{1}(\mathbb{R}^{N})$ be a solution of (2.1). Setting $a(x)=\varepsilon^{2}+|u(x)|^{p-2}$ we see that $|-\varepsilon^{2}u+Q(x)|u|^{p-2}u|\leq a(x)|u|$ and $a\in L^{\frac{N}{2}}_{loc}(\mathbb{R}^{N})$. Then, a Brezis-Kato-Moser argument shows that $u\in L^{s}_{loc}(\mathbb{R}^{N})$ for every $s\in[1,\infty)$ (see [14, Lemma B.3]), and standard elliptic regularity arguments yield that $u\in\mathcal{C}^{1,\alpha}_{loc}(\mathbb{R}^{N})$ for every $\alpha\in(0,1)$ and $u\in\mathcal{C}_{loc}^{\infty}(\Omega)\cap\mathcal{C}_{loc}^{\infty}(\mathbb{R}^{N}\smallsetminus\overline{\Omega})$ (see [14, Appendix B]). ∎
@@
The existence of a least energy solution follows from Ekeland’s variational principle and Lemma 2.3. Replacing it by its absolute value, we obtain a least energy solution such that $u\geq 0$ in $\mathbb{R}^{N}$. As shown in Lemma 2.4, $u\in W^{2,N}_{loc}(\mathbb{R}^{N})$. Thus, the strong maximum principle for strong solutions [9, Theorem 9.6] applied to the equation
@@
is continuous and odd. Therefore, $\infty=\mathrm{genus}(\Sigma_{\Omega})\leq\mathrm{genus}(\mathcal{N}_{\varepsilon})$, where “$\mathrm{genus}$” stands for the Kransnoselskii genus. It follows from [14, Theorem II.5.7] that $J_{\varepsilon}$ has infinitely many pairs of critical points on $\mathcal{N}_{\varepsilon}$ and following the argument in [12, Proposition 9.33] one shows that $J_{\varepsilon}$ has an unbounded sequence of critical values on $\mathcal{N}_{\varepsilon}$, as claimed. ∎
@@ 3 The limit profile of positive least energy solutions
When $\Omega$ is the unit ball, the behavior of positive least energy solutions to (2.1) as $\varepsilon\to 0$ is described in [8, Theorem 1.1]. It is shown that, after passing to a subsequence they converge strongly in $E$ to a positive least energy solution to the problem (1.2). The general case is similar. We give the details.
@@
that $u_{k}\to u$ strongly in $D^{1,2}(\mathbb{R}^{N})$ and that $t=1$. Hence, $\|u\|^{2}=\int_{\mathbb{R}^{N}}Q(x)|u|^{p}=\frac{2p}{p-2}c_{0}$. Combining this identity with (3.4) we see that $u_{k}\to u$ strongly in $L^{p}(\mathbb{R}^{N})$ and, thus, in $E$. As $u\in\mathcal{N}_{0}$ and $J_{0}(u)=c_{0}$, $u$ is a least energy solution to (1.2). Since $u\geq 0$, from Lemma 3.3 and the maximum principle (see, for instance, [9, Theorem 9.6]) we derive that $u>0$ in $\mathbb{R}^{N}$. ∎
@@ 4 Existence of least energy nodal solutions
$\gamma(s)\in\widehat{\mathcal{N}}_{\varepsilon}^{+}\text{ \ iff \ }s\in(0,\tfrac{1}{2})\qquad\text{and}\qquad\gamma(s)\in\widehat{\mathcal{N}}_{\varepsilon}^{-}\text{ \ iff \ }s\in(\tfrac{1}{2},1).$
@@
Given $u\in\mathcal{E}_{\varepsilon}$, define $\gamma=\gamma_{u}:[0,1]\to\mathcal{N}_{\varepsilon}$ as
$\gamma(s):=t_{u_{s}}u_{s},\qquad\text{where}\quad u_{s}:=(1-s)u^{+}+su^{-}$
and $t_{u_{s}}$ is the unique positive number such that $t_{u_{s}}u_{s}\in\mathcal{N}_{\varepsilon}$, as in Lemma 2.1. Note that $u_{s}\in\mathcal{U}$ for every $s\in[0,1]$, so $t_{u_{s}}$ is well defined. As $u^{+},u^{-},u\in\mathcal{N}_{\varepsilon}$, this path satisfies
$\gamma(0)=u^{+},\qquad\gamma(1)=u^{-},\qquad\gamma\left(\tfrac{1}{2}\right)=u.$
Fix $s\in(0,1)$ and set $w:=\gamma(s)$. Then, $w^{+}=t_{u_{s}}(1-s)u^{+}$ and $w^{-}=t_{u_{s}}su^{-}$. Since $w\in\mathcal{N}_{\varepsilon}$ we have that
$\|w^{+}\|^{2}+\|w^{-}\|^{2}=\|w\|^{2}=\int_{\mathbb{R}^{N}}Q(x)|w|^{p}=\int_{\mathbb{R}^{N}}Q(x)|w^{+}|^{p}+\int_{\mathbb{R}^{N}}Q(x)|w^{-}|^{p}.$
Therefore, recalling that $u^{+},u^{-}\in\mathcal{N}_{\varepsilon}$, we obtain
$w\in\widehat{\mathcal{N}}_{\varepsilon}^{+}$ $\quad\Longleftrightarrow\quad\|w^{+}\|^{2}<\int_{\mathbb{R}^{N}}Q(x)|w^{+}|^{p}\quad\Longleftrightarrow\quad\|w^{-}\|^{2}>\int_{\mathbb{R}^{N}}Q(x)|w^{-}|^{p}$
$\quad\Longleftrightarrow\quad(t_{u_{s}}(1-s))^{p-2}>1>(t_{u_{s}}s)^{p-2}\quad\Longleftrightarrow\quad s\in(0,\tfrac{1}{2}).$
Similarly,
$w\in\widehat{\mathcal{N}}_{\varepsilon}^{-}\quad\Longleftrightarrow\quad s\in(\tfrac{1}{2},1).$
Next, fix $s\in(0,\frac{1}{2})$. Note that $t_{w^{+}}=(t_{u_{s}}(1-s))^{-1}$. Then, from Lemma 2.1, we deduce that $J_{\varepsilon}(t_{w^{+}}w^{-})=J_{\varepsilon}(\frac{s}{1-s}u^{-})>0$ and
@@
$<J_{\varepsilon}(w)=J_{\varepsilon}(t_{u_{s}}(1-s)u^{+})+J_{\varepsilon}(t_{u_{s}}su^{-})<J_{\varepsilon}(u^{+})+J_{\varepsilon}(u^{-})=J_{\varepsilon}(u),$
as claimed. The argument for $s\in(\frac{1}{2},1)$ is similar. This completes the proof. ∎
@@
Arguing by contradiction, assume that $J_{\varepsilon}^{\prime}(u)\neq 0$. Fix $\delta>0$ such that $\overline{B}_{\delta}(u):=\{v\in\mathcal{N}_{\varepsilon}:\|v-u\|_{\varepsilon}\leq\delta\}\subset\widehat{\mathcal{N}}_{\varepsilon}$ and $\|\nabla_{\mathcal{N}_{\varepsilon}}J_{\varepsilon}(v)\|\geq\alpha>0$ for every $v\in\overline{B}_{\delta}(u)$, where $\nabla_{\mathcal{N}_{\varepsilon}}J_{\varepsilon}(v)$ is the gradient of $J_{\varepsilon}(v)|_{\mathcal{N}_{\varepsilon}}$ at $v$. Choose $\beta\in(0,\frac{\delta\alpha}{16})$. The deformation lemma [19, Lemma 5.15] yields a map $\eta:[0,1]\times\mathcal{N}_{\varepsilon}\to\mathcal{N}_{\varepsilon}$ such that $\eta(t,u)=u$ if $t=0$ or $u\notin J_{\varepsilon}^{-1}[\widehat{c}_{\varepsilon}-2\beta,\widehat{c}_{\varepsilon}+2\beta]\cap\overline{B}_{\delta}(u)$, $J_{\varepsilon}(\eta(1,u))\leq\widehat{c}_{\varepsilon}-\beta$ and $t\mapsto J_{\varepsilon}(\eta(t,v))$ is nonincreasing for every $v\in\mathcal{N}_{\varepsilon}$. Furthermore, $\eta(t,v)\in\overline{B}_{\delta}(u)$ if $v\in\overline{B}_{\delta}(u)$ for every $t\in[0,1]$. Define $\sigma:[0,1]\to\mathcal{N}_{\varepsilon}$ by $\sigma(s):=\eta(1,\gamma(s))$. Then,
@@ 6.1 Radial symmetry of the positive least energy solution
We recall the notion of Schwarz symmetrization, also called symmetric rearrangement (see [10, Chapter 1] or [11]). Let $A$ be a measurable set of finite volume $|A|$ in $\mathbb{R}^{N}$. Its Schwarz symmetrization $A^{*}$ is the open ball in $\mathbb{R}^{N}$ centered at the origin whose volume is the same as the volume of $A$. Let $u:\mathbb{R}^{N}\to\mathbb{R}$ be a nonnegative measurable function that vanishes at infinity, in the sense that its level sets have finite volume, i.e.,
@@
These and other properties may be found in [11, Section 3.3].
@@
Let $w$ be a positive least energy solution of (1.2) with $\Omega=B_{1}$. By [11, Lemma 7.17] we have that $\|w\|^{2}\geq\|w^{*}\|^{2}$ and, as pointed out above, $|w|_{p}=|w^{*}|_{p}$. We claim that
@@ 6.2 Foliated Schwarz symmetry of the least energy nodal solution
Let $\mathbb{S}^{N-1}=\left\{x\in\mathbb{R}^{N}:|x|=1\right\}$ be the unit sphere. For each $e\in\mathbb{S}^{N-1}$ we consider the halfspace $H(e):=\{x\in$ $\left.\mathbb{R}^{N}:x\cdot e>0\right\}$, and we denote by $u_{e}$ the composition of the function $u:\mathbb{R}^{N}\rightarrow\mathbb{R}$ with the reflection $\sigma_{e}$ with respect to the hyperplane $\partial H(e)$, that is,
@@
Observe that, for any $u:\mathbb{R}^{N}\to\mathbb{R}$, we have that $(u^{\pm})_{H}=(u_{H})^{\pm}$ and that $(u\chi_{A})_{H}=u_{H}\chi_{A}$ for any radially symmetric set $A$. From [18, Corollary 3.2] we have that
@@
and, from [18, Lemma 3.1], we derive
@@
Let $\Omega$ be a radially symmetric open set in $\mathbb{R}^{N}$. There exists $e^{*}\in\mathbb{S}^{N-1}$ such that $u\in\mathcal{C}^{0}(\overline{\Omega})$ is foliated Schwarz symmetric with respect to $e^{*}$ if and only if for every $e\in\mathbb{S}^{N-1}$ either
@@
The proof is given in [3] (see also [13, Proposition 3.2] and [18]). This proof only takes into account the case when $\Omega$ is connected, but the general case easily follows since every component satisfies the same inequalities which completely determine the symmetry axis and the angular monotonicity. ∎
@@
Clearly, $|u-u_{e}|=0$ on $\partial H(e)$. By the maximum principle for strong solutions [9, Theorem 9.6], either $u\equiv u_{e}$ in $H(e)\cap\Omega$ or $\left|u-u_{e}\right|>0$ in $H(e)\cap\Omega$. Since $e$ is arbitrary, Lemma 6.3 states that there exists $e^{*}\in\mathbb{S}^{N-1}$ such that $u$ is foliated Schwarz symmetric in $\Omega$ with respect to $e^{*}$. In particular, it is axially symmetric in $\Omega$ with respect to $\mathbb{R}e^{*}$. Therefore,